### v1

@@ 2.3 Empirical Corpus
Given the diverse nature of toxicity, ranging from offensive language to harmful behavior, datasets must be tailored to the specific type of content they aim to analyze. The purpose of this aggregation is to provide a comprehensive overview of the datasets used in toxicity research, with a focus on both SE and non-SE contexts. By distinguishing between SE and non-SE datasets, we can better understand the unique challenges and characteristics of toxicity in different environments. SE datasets, often derived from developer communication platforms such as code reviews, issue trackers, and forums, present unique challenges such as the need to account for technical jargon and collaborative discourse. On the other hand, non-SE datasets, which may include social media, news articles, or general-purpose comment sections, often feature broader forms of toxicity, including hate speech, cyberbullying, and harassment. This segmentation allows researchers to identify common trends, evaluate the applicability of models across domains, and highlight areas where further research and dataset development are needed.
Tables 4 and 5 provide a summary of key SE and non-SE datasets used in toxicity detection research, organized by their characteristics, sources, annotation methods, and evaluation metrics. This overview provides a foundational reference for researchers seeking to build on existing work, compare approaches, and identify gaps in dataset development to enhance toxicity detection systems.
@@ 2.4 Study Selection
To identify relevant articles, we conducted a systematic search across four major academic databases: IEEE Xplore, ACM Digital Library, Springer, and arXiv. Given our focus on recent advancements in toxicity research within software engineering, we restricted our search to articles published between 2020 and 2025. Our initial query included the keywords toxicity and software engineering to retrieve a broad yet relevant set of studies.
The selection process consisted of four sequential steps. First, we screened titles, abstracts, and keywords to ensure the studies were relevant, scientifically rigorous, and offered meaningful insights. Second, we assessed the credibility of each study by evaluating the publication venues. Third, we removed duplicate entries to maintain a unique dataset. Finally, we conducted a thorough qualitative assessment to identify the most relevant and influential studies. After selecting these core studies, we applied a snowballing approach, tracing citations both forward and backward, ultimately identifying a total of 63 papers for inclusion in this survey. The complete process is illustrated in Figure 2.
Figure 2: Study Identification and Selection Process
Table 3: Overview of Core Literature on Toxicity Study in Software Engineering
@@
1 ToxiSpanSE: An explainable toxicity detection in code review comments ESEM  2023 SE datasets, Human annotator, Span-based detector, Domain-specific tailoring, LLM based model, Feature engineering Yes No
2 A benchmark study of the contemporary toxicity detectors on software engineering interactions APSEC  2020 SE datasets, Non-SE datasets, Human annotator, Algorithmic annotator, Domain-specific tailoring Yes No
3 Exploring ChatGPT for Toxicity Detection in GitHub ICSE  2024 SE datasets, LLM based model, Domain-specific tailoring, Human annotator Yes No
4 Stress and burnout in open source: Toward finding, understanding, and mitigating unhealthy interactions ICSE  2020 SE datasets, Domain-specific tailoring, Human annotator, Feature engineering Yes No
5 An Empirical Evaluation of the Zero-shot, Few-shot, and Traditional Fine-tuning Based Pretrained Language Models for Sentiment Analysis in Software Engineering IEEE Access  2024 SE datasets, Human annotator, Feature engineering, LLM based model, Domain-specific tailoring Yes No
6 Identification and mitigation of toxic communications among open source software developers ASE  2022a SE datasets, Human annotator, Data cleaning, Feature engineering, Domain-specific tailoring Yes No
7 Towards offensive language detection and reduction in four software engineering communities EASE  2021 SE Datasets, Human annotator, Feature engineering, Domain-specific tailoring Yes Yes
8 Realtoxicityprompts: Evaluating neural toxic degeneration in language models arXiv preprint  2020 Non-SE datasets, Algorithmic annotator, Data cleaning, LLM based model Yes No
9 Exploring ChatGPT for identifying sexism in the communication of software developers PETRA  2024 LLM based model, Prompt engineering Yes No
10 DeMod: A Holistic Tool with Explainable Detection and Personalized Modification for Toxicity Censorship arXiv preprint  2024 Non-SE Datasets, Human annotator, LLM based model, Domain-specific tailoring, Span-based detector Yes Yes
@@
ToxiCR Dataset Toxicity Detection FOSS project code reviews and Gitter chat messages https://github.com/WSU-SEAL/ToxiCR 20k Manual labeling (2 of the authors)  2022a ASE Accuracy, Precision, Recall, F-score, Cohen’s $\kappa$
Benchmark Study Dataset Toxicity Detection FOSS project code reviews, Gitter chat messages, and Jigsaw test sample https://github.com/WSU-SEAL/toxicity-dataset/ 14k Manual labeling (2 of the authors)  2020 APSEC Accuracy, Precision, Recall, F-score, Cohen’s $\kappa$
ToxiSpanSE Toxic Span Detection ToxiCR Dataset https://github.com/WSU-SEAL/ToxiSpanSE 23k Manual labeling (2 independent annotators)  2023 ESEM Krippendorff’s $\alpha$, F1
GitHub Issues Dataset Toxicity Detection GitHub https://github.com/CMUSTRUDEL/toxicity-detector 2k Perspective API  2020 ICSE F0.5
Comment Dataset Toxicity Detection GitHub, Gitter, Slack, Stack Overflow https://archive.org/details/stackexchange 1,130k Perspective API and Regex  2021 EASE Precision
@@
Real Toxicity Prompts Toxicity Detection Reddit https://github.com/allenai/real-toxicity-prompts 22k Perspective API  2020 EMNLP Expected Maximum Toxicity and Toxicity Probability
Writing Prompts Toxicity Mitigation Reddit https://www.kaggle.com/datasets/ratthachat/writing-prompts 300k Toxicity detection API  2022 AAAI Toxicity Reduction
Offensive Language Identification Dataset (OLID) Offensive Language Detection Twitter https://sites.google.com/site/offensevalsharedtask/olid 14.2k 3 annotators per rater group per comment  2019 NAACL Per-class Precision, Recall, F1, Weighted Average
Hate Speech Dataset Hate Speech, Offensive Language Detection Twitter https://github.com/BenchengW/Offensive-Language-Detection-DL 25k 3 to 9 annotators per post  2020 AI2SD Accuracy
@@ 3 Methodology
This section outlines common methodologies for addressing toxicity in SE. We first discuss annotation techniques involving human and algorithmic annotators. Next, we describe preprocessing methods, including data cleaning, feature engineering and sampling techniques. Finally, we discussed various models from multiple perspectives, including general versus domain-specific, model architecture and task types.
@@ 3.1.1 Human annotator
The involvement of human annotators is indispensable for capturing context-specific nuances in toxic content. To enhance annotation quality, researchers employ several strategies. One widely adopted approach is the use of multiple annotators with diverse backgrounds, which helps mitigate individual biases and improve perspective diversity. For instance, Sarker et al. [47] paired annotators of different genders to label samples, increasing the likelihood of balanced viewpoint representation.
@@
To ensure consistent labeling, annotators often follow a well-defined rubric that outlines toxicity criteria. These rubrics can be pre-established or developed specifically for a given research context. For example, in an effort to tailor annotation guidelines to the SE domain, Sarker et al. [46] had a group of annotators analyze 1000 SE texts and empirically develop a rubric for identifying toxicity. This process not only clarified labeling standards but also reduced subjectivity by aligning annotators with detailed, research-specific rules. The resulting rubric was later reused in subsequent work [44], demonstrating its reliability and applicability across studies.
@@
Beyond following structured guidelines, researchers also assess annotation quality through inter-annotator agreement metrics such as Cohen’s $\kappa$ or Krippendorff’s $\alpha$. The choice of metric depends on the complexity of the task and the specific research needs. For instance, Sarker et al. [47] utilized Krippendorff’s $\alpha$, which accounts for multiple annotators, missing values, and partial agreements—making it particularly suitable for nuanced tasks like toxicity annotation. By quantifying consistency between annotators, these metrics help validate dataset reliability. When disagreements arise, conflict resolution mechanisms are employed to establish a reliable ground truth. Typically, a third annotator or a consensus process determines the final label, ensuring that conflicting perspectives are reconciled.
@@
Additionally, multi-label annotation offers a more nuanced alternative to traditional binary classification, which often oversimplifies the complexity of toxic language. Instead of strictly categorizing content as offensive or non-offensive, multi-label annotation captures contextual dependencies and annotators’ varying perspectives.  Arhin et al. [5] introduced an approach incorporating three label types—normal, offensive, and undecided—along with three labeling schemes (strict, relaxed, and inferred group labels). This method enables a more comprehensive and context-aware annotation process, ultimately leading to more robust toxicity detection models.
@@ 3.1.2 Algorithmic Annotator
To reduce the costs and time associated with manual annotation, algorithmic tools are increasingly employed in toxicity research. These tools enable efficient labeling of large-scale corpora, providing automated assessments that assist researchers in filtering and prioritizing data for analysis.
@@
One widely used tool in this domain is the Google Perspective API [14], which assigns toxicity scores to text based on linguistic attributes such as ”toxicity,” ”insult,” and ”identity attack.” Developed by Google Jigsaw, this API leverages machine learning models trained on extensive datasets to identify harmful or offensive language patterns in online communication. The Perspective API is particularly useful for pre-labeling datasets, allowing researchers to streamline manual annotation efforts while gaining a granular understanding of toxicity intensity. Its scoring system facilitates the evaluation of language models and informs the development of mitigation strategies. Given its effectiveness, studies [13, 20, 41] frequently use Perspective API as a benchmarking tool for validating toxicity detection methods. Additionally, it is often integrated into hybrid annotation pipelines that combine algorithmic and human inputs, thereby improving data quality and model robustness.
@@
However, despite its widespread adoption, the Perspective API has notable limitations and potential biases that impact its reliability and fairness. Studies have highlighted systematic biases in its scoring, particularly against marginalized groups. For example, Sap et al. [43] demonstrated that the API disproportionately flags African American Vernacular English (AAVE) as more toxic than standard English, underscoring the risks of algorithmic discrimination in automated moderation tools. Furthermore, the API’s lack of transparency in its training data and model updates presents challenges for reproducibility and fairness auditing. Bender et al. [9] emphasize that black-box models like Perspective API hinder researchers’ ability to critically assess how biases emerge, making it difficult to address or mitigate them. Without open access to the underlying data, researchers must rely on empirical testing to infer weaknesses, which limits their ability to propose targeted improvements.
@@
In addition to machine learning-based approaches, researchers also employ simpler rule-based methods such as regular expression (regex) matching to detect explicit toxic content. Regex allows for the definition of patterns that capture specific linguistic constructs, including slurs, profanities, or abusive language, enabling rapid identification of harmful text. For instance, regex can be used to flag messages containing offensive words, including obfuscated spellings such as ”f**k.” The primary advantage of regex-based methods lies in their speed and cost-effectiveness, as they require minimal computational resources compared to more sophisticated machine learning models or APIs. However, regex-based approaches have significant limitations: they struggle to interpret context, irony, or subtle variations in language, often leading to high false positive or negative rates. Nozza et al. [37] discussed the complexities introduced by various obfuscation methods, such as character substitutions, which complicate the detection process and require regex patterns to account for numerous variations. For example, crafting regex patterns to identify variations like ”f**k” requires anticipating and encoding potential character substitutions. Despite these drawbacks, regex remains a valuable tool for filtering explicit toxicity, particularly when used in conjunction with more advanced annotation techniques.
3.2 Preprocessing

### v2

@@ 2.3 Empirical corpus
Given toxicity’s diverse nature, from offensive language to harmful behavior, datasets must be specifically tailored. This section provides a comprehensive overview of toxicity datasets, focusing on both SE and non-SE contexts. Distinguishing these domains is crucial for understanding their unique challenges and characteristics: SE datasets such as raman2020stress and sarker2023toxispanse, from developer communication platforms, present challenges like technical jargon and collaborative conversations, while non-SE datasets such as gehman2020realtoxicityprompts and xu2022leashing often feature broader forms of toxicity, including hate speech and cyberbullying. This segmentation allows researchers to identify common trends, evaluate cross-domain model applicability, and highlight areas for future dataset development.
However, existing datasets, particularly in the SE domain, often present notable limitations regarding representativeness and sampling biases. A pervasive issue is significant class imbalance, with toxic instances being rare; for example, the ToxiCR dataset contains only 4.4% toxic instances [sarker2023automated]. This extreme rarity makes it unfeasible to gather sufficient toxic examples via random sampling. Consequently, researchers must employ specific, often biased, sampling strategies. While efforts like combining multiple detection methods (e.g., language-based detectors, locked issues) to enhance sample diversity [miller2022did] are crucial for data collection, they inherently risk biasing the type of toxicity captured. This known trade-off between sample size and representativeness limits the broad applicability of resulting datasets and can cause models to struggle in effectively detecting the minority (toxic) class, thereby restricting their practical utility. This dataset imbalance can be addressed in future work by leveraging over-sampling or down-sampling methods to synthesize minority labels, such as SMOTE (Synthetic Minority Over-sampling Technique) [chawla2002smote], which creates synthetic examples for the minority class by interpolating between existing minority instances rather than simply duplicating them.
Furthermore, current SE toxicity datasets are overwhelmingly English-centric, an oversight given the global, multilingual nature of software development. While multilingual data have been utilized to study toxicity in general online discourse [leite2020toxiclanguagedetectionsocial], its application in SE has been less prominent. This English predominance, with limited non-English SE contexts (e.g., Chinese or Arabic developer forums), suggests a linguistic bias potentially affecting research generalizability. Thus, curating multilingual SE toxicity data is crucial for broader coverage and more inclusive detection systems worldwide.
Tables 6 and 7 provide a summary of key SE and non-SE datasets used in toxicity detection research, organized by their characteristics, sources, annotation methods, and evaluation metrics. This overview provides a foundational reference for researchers seeking to build on existing work, compare approaches, and identify gaps in dataset development to enhance toxicity detection systems.
@@ 2.4 Study selection
To identify relevant articles, we conducted a systematic search across four major academic databases: IEEE Xplore, ACM Digital Library, Springer, and the preprint server, arXiv. The inclusion of arXiv is crucial for capturing the most recent, innovative, and early-stage findings in the dynamic LLM-toxicity intersection before formal peer review. This practice aligns with best practices for multivocal literature reviews (MLRs) in software engineering, where incorporating grey literature like preprints is recommended to capture current knowledge [garousi2019guidelines]. The risk associated with including unreviewed preprints is mitigated by a subsequent rigorous screening process (detailed below). Given our focus on recent software engineering advancements in toxicity, we restricted our search to articles published between 2020 and 2024.
Our initial search query, which utilized the core terms toxic and large language models, yielded an initial pool of 2,280 studies. The selection of these keywords was systematically motivated by their effectiveness in maximizing recall. We performed initial pilot searches to scope the vocabulary used in this domain, specifically comparing the noun toxicity with the adjective toxic. The latter provided significantly broader coverage in the target databases and was thus selected as the primary identifier for this dimension of the search to strategically prioritize recall over precision. The specific queries used for each of the libraries are detailed in Table 3.
The selection of primary studies followed a structured, multi-stage screening method to ensure the relevance, quality, and uniqueness of the final study set, starting from an initial pool of 2,280 studies. This general process adheres to the well-established guidelines for conducting Systematic Literature Reviews (SLRs) in SE [kitchenham2009systematic].
Studies were included based on the following criteria:
  • I1: Credibility: Publication in highly-ranked conferences or journals in the SE and NLP domains, as determined by a venue assessment.
  • I2: Core Focus and Methodology: Must be an empirical study presenting concrete results related to toxicity detection or mitigation strategies in SE, specifically utilizing or developing LLM methods and relying on human-annotated ground truth.
  • I3: Data Quality: For studies utilizing human-annotated data, inter-annotator agreement must be at least $0.6$ for Cohen’s $\kappa$ [cohen1960coefficient] or $0.8$ for Krippendorff’s $\alpha$ [krippendorff2018content].
Studies were excluded based on the following criteria:
  • E1: Topic Mismatch: Purely conceptual works, or studies not directly focused on toxicity within SE.
  • E2: Language Mismatch: Non-English language papers.
  • E3: Duplicates: Identification as a duplicate entry across databases.
This rigorous, four-step process is outlined in Table  4, resulting in a final corpus of 63 unique and highly relevant primary studies for this survey. The complete process is illustrated in Figure 2.
Table 3: Queries to four digital libraries to identify relevant works
Library Query String Filters Result
IEEE Xplore (“All Metadata”:toxic large language models) Conferences; 2020-2024 40
ACM [All: toxic large language models] AND [E-Publication Date: (01/01/2020 TO 12/31/2024)] Research Article 1480
Springer toxic large language models) Conference paper; 2020-2024 437
arXiv order: -announced_date_first; size: 50; date_range: from 2020-01-01 to 2024-12-31; classification: Computer Science (cs); include_cross_list: True; terms: AND all=toxic large language models N/A 323
Table 4: Summary of the four-step study selection process
Step Process Description Initial Pool Excluded Resulting Pool
1 Preliminary Screening: Title, abstract, and keyword screening (E1, E2) was performed to ensure basic relevance, language, and publication time. 2,280 2,043 237
2 Credibility Assessment: A full-text evaluation and publication venue check was performed against criterion I1, prioritizing highly-ranked venues. 237 166 71
3 Duplicate Removal: A check for unique entries was conducted against criterion E3. 71 19 52
4 Qualitative Assessment & Snowballing: The 52 unique studies underwent a thorough qualitative assessment against criteria I2 and I3, focusing on LLM methods and data quality, resulting in 10 core studies, as detailed in Table 5. A subsequent snowballing approach was applied to this core literature. 52 42 63 (after snowballing)
Figure 2: Study identification and selection process applied in this literature review
Table 5: Overview of core literature on toxicity study in software engineering.
@@
1 ToxiSpanSE: An explainable toxicity detection in code review comments ESEM  sarker2023toxispanse SE datasets, Human annotator, Span-based detector, Domain-specific tailoring, LLM based model, Feature engineering Yes No
2 A benchmark study of the contemporary toxicity detectors on software engineering interactions APSEC  sarker2020benchmark SE datasets, Non-SE datasets, Human annotator, Algorithmic annotator, Domain-specific tailoring Yes No
3 Exploring ChatGPT for Toxicity Detection in GitHub ICSE  mishra2024exploring SE datasets, LLM based model, Domain-specific tailoring, Human annotator Yes No
4 Stress and burnout in open source: Toward finding, understanding, and mitigating unhealthy interactions ICSE  raman2020stress SE datasets, Domain-specific tailoring, Human annotator, Feature engineering Yes No
5 An Empirical Evaluation of the Zero-shot, Few-shot, and Traditional Fine-tuning Based Pretrained Language Models for Sentiment Analysis in Software Engineering IEEE Access  shafikuzzaman2024empirical SE datasets, Human annotator, Feature engineering, LLM based model, Domain-specific tailoring Yes No
6 Identification and mitigation of toxic communications among open source software developers ASE  sarker2022identification SE datasets, Human annotator, Data cleaning, Feature engineering, Domain-specific tailoring Yes No
7 Towards offensive language detection and reduction in four software engineering communities EASE  cheriyan2021towards SE Datasets, Human annotator, Feature engineering, Domain-specific tailoring Yes Yes
8 Realtoxicityprompts: Evaluating neural toxic degeneration in language models arXiv preprint  gehman2020realtoxicityprompts Non-SE datasets, Algorithmic annotator, Data cleaning, LLM based model Yes No
9 Exploring ChatGPT for identifying sexism in the communication of software developers PETRA  sultana2024exploring LLM based model, Prompt engineering Yes No
10 DeMod: A Holistic Tool with Explainable Detection and Personalized Modification for Toxicity Censorship arXiv preprint  li2024demod Non-SE Datasets, Human annotator, LLM based model, Domain-specific tailoring, Span-based detector Yes Yes
@@
ToxiCR Dataset Toxicity Detection FOSS project code reviews and Gitter chat messages https://github.com/WSU-SEAL/ToxiCR 20k Manual labeling (2 of the authors)  sarker2022identification ASE Accuracy, Precision, Recall, F-score, Cohen’s $\kappa$
Benchmark Study Dataset Toxicity Detection FOSS project code reviews, Gitter chat messages, and Jigsaw test sample https://github.com/WSU-SEAL/toxicity-dataset/ 14k Manual labeling (2 of the authors)  sarker2020benchmark APSEC Accuracy, Precision, Recall, F-score, Cohen’s $\kappa$
ToxiSpanSE Toxic Span Detection ToxiCR Dataset https://github.com/WSU-SEAL/ToxiSpanSE 23k Manual labeling (2 independent annotators)  sarker2023toxispanse ESEM Krippendorff’s $\alpha$, F1
GitHub Issues Dataset Toxicity Detection GitHub https://github.com/CMUSTRUDEL/toxicity-detector 2k Perspective API  raman2020stress ICSE F0.5
Comment Dataset Toxicity Detection GitHub, Gitter, Slack, Stack Overflow https://archive.org/details/stackexchange 1,130k Perspective API and Regex  cheriyan2021towards EASE Precision
@@
Real Toxicity Prompts Toxicity Detection Reddit https://github.com/allenai/real-toxicity-prompts 22k Perspective API  gehman2020realtoxicityprompts EMNLP Expected Maximum Toxicity and Toxicity Probability
Writing Prompts Toxicity Mitigation Reddit https://www.kaggle.com/datasets/ratthachat/writing-prompts 300k Toxicity detection API  xu2022leashing AAAI Toxicity Reduction
Offensive Language Identification Dataset (OLID) Offensive Language Detection Twitter https://sites.google.com/site/offensevalsharedtask/olid 14.2k 3 annotators per rater group per comment  zampieri2019predicting NAACL Per-class Precision, Recall, F1, Weighted Average
Hate Speech Dataset Hate Speech, Offensive Language Detection Twitter https://github.com/BenchengW/Offensive-Language-Detection-DL 25k 3 to 9 annotators per post  touahri2020offensive AI2SD Accuracy
@@ 3 Methodology
This section outlines common methodologies for addressing toxicity in SE. We first discuss annotation techniques involving human and algorithmic annotators. Next, we describe pre-processing methods, including data cleaning, feature engineering and sampling techniques. Finally, we discussed various models from multiple perspectives, including general versus domain-specific, model architecture, and task types.
@@ 3.1.1 Human annotator
The involvement of human annotators is indispensable for capturing context-specific nuances in toxic content. To enhance annotation quality, researchers employ several strategies. One widely adopted approach is the use of multiple annotators with diverse backgrounds, which helps mitigate individual biases and improve perspective diversity. For instance, sarker2023toxispanse paired annotators of different genders to label samples, increasing the likelihood of balanced viewpoint representation.
@@
To ensure consistent labeling, annotators often follow a well-defined rubric that outlines toxicity criteria. These rubrics can be pre-established or developed specifically for a given research context. For example, in an effort to tailor annotation guidelines to the SE domain, sarker2020benchmark had a group of annotators analyze 1000 SE texts and empirically develop a rubric for identifying toxicity. This process not only clarified labeling standards but also reduced subjectivity by aligning annotators with detailed, research-specific rules. The resulting rubric was later reused in subsequent work [sarker2022identification], demonstrating its reliability and applicability across studies.
@@
Beyond following structured guidelines, researchers also assess annotation quality through inter-annotator agreement metrics such as Cohen’s $\kappa$ or Krippendorff’s $\alpha$. The choice of metric depends on the complexity of the task and the specific research needs. For instance, sarker2023toxispanse utilized Krippendorff’s $\alpha$, which accounts for multiple annotators, missing values, and partial agreements—making it particularly suitable for nuanced tasks like toxicity annotation. By quantifying consistency between annotators, these metrics help validate dataset reliability. When disagreements arise, conflict resolution mechanisms are employed to establish a reliable ground truth. Typically, a third annotator or a consensus process determines the final label, ensuring that conflicting perspectives are reconciled.
For practical implementation in annotation teams, particularly within the SE domain, establishing clear protocols for annotator training, target inter-annotator agreement (IAA), and conflict resolution is crucial. Training protocols often involve initial calibration sessions where annotators discuss guidelines and ambiguities, ensuring a shared understanding of criteria. While specific benchmarks for IAA can vary, aiming for high agreement is essential for data reliability. For instance, Cohen’s $\kappa$ is typically considered to indicate substantial agreement above 0.7, and studies like sarker2023toxispanse have reported agreement as high as 0.96. When disagreements arise, an adjudication workflow, such as facilitated discussions or involvement of a senior annotator, is necessary to reach consensus and establish a robust ground truth for conflicting labels.
@@
Additionally, multi-label annotation offers a more nuanced alternative to traditional binary classification, which often oversimplifies the complexity of toxic language. Instead of strictly categorizing content as offensive or non-offensive, multi-label annotation captures contextual dependencies and annotators’ varying perspectives.  arhin2021ground introduced an approach incorporating three label types—normal, offensive, and undecided—along with three labeling schemes (strict, relaxed, and inferred group labels). This method enables a more comprehensive and context-aware annotation process, ultimately leading to more robust toxicity detection models.
@@ 3.1.2 Algorithmic annotator
To reduce the costs and time associated with manual annotation, algorithmic tools are increasingly employed in toxicity research. These tools enable efficient labeling of large-scale corpora, providing automated assessments that assist researchers in filtering and prioritizing data for analysis. Among these, several commercial APIs and open-source libraries offer automated assessment capabilities.
One widely used tool in this domain is the Google Perspective API [conversationAI], which assigns toxicity scores to text based on linguistic attributes such as “toxicity”, “insult,“ and “identity attack”. Developed by Google Jigsaw, this API leverages machine learning models trained on extensive datasets to identify harmful or offensive language patterns in online communication. Beyond Perspective API, other notable commercial services include IBM Watson’s Tone Analyzer, which offers sentiment and tone analysis for various communication styles [tone2018], and Microsoft’s Azure AI Content Safety (formerly Content Moderator), designed to detect objectionable content across text and images [azure2025]. Researchers also widely utilize open-source libraries such as Detoxify, which provides flexible and customizable solutions for toxicity classification [detoxify2021].
@@
The Perspective API is particularly useful for pre-labeling datasets, allowing researchers to streamline manual annotation efforts while gaining a granular understanding of toxicity intensity. Its scoring system facilitates the evaluation of language models and informs the development of mitigation strategies. Given its effectiveness, studies [cheriyan2021towards, gehman2020realtoxicityprompts, raman2020stress] frequently use Perspective API as a benchmarking tool for validating toxicity detection methods. Additionally, it is often integrated into hybrid annotation pipelines that combine algorithmic and human inputs, thereby improving data quality and model robustness.
@@
However, despite its widespread adoption, the Perspective API has notable limitations and potential biases that impact its reliability and fairness. Studies have highlighted systematic biases in its scoring, particularly against marginalized groups. For example, sap2019risk demonstrated that the API disproportionately flags African American Vernacular English (AAVE) as more toxic than standard English, underscoring the risks of algorithmic discrimination in automated moderation tools. Furthermore, the API’s lack of transparency in its training data and model updates presents challenges for reproducibility and fairness auditing. bender2021dangers emphasize that black-box models like Perspective API hinder researchers’ ability to critically assess how biases emerge, making it difficult to address or mitigate them. Without open access to the underlying data, researchers must rely on empirical testing to infer weaknesses, which limits their ability to propose targeted improvements.
@@
In addition to machine learning-based approaches, researchers also employ simpler rule-based methods such as regular expression (regex) matching to detect explicit toxic content. Regex allows for the definition of patterns that capture specific linguistic constructs, including slurs, profanities, or abusive language, enabling rapid identification of harmful text. For instance, regex can be used to flag messages containing offensive words, including obfuscated spellings such as “f**k.” The primary advantage of regex-based methods lies in their speed and cost-effectiveness, as they require minimal computational resources compared to more sophisticated machine learning models or APIs. However, regex-based approaches have significant limitations: they struggle to interpret context, irony, or subtle variations in language, often leading to high false positive or negative rates. nozza2023state discussed the complexities introduced by various obfuscation methods, such as character substitutions, which complicate the detection process and require regex patterns to account for numerous variations. For example, crafting regex patterns to identify variations like “f**k” requires anticipating and encoding potential character substitutions. Despite these drawbacks, regex remains a valuable tool for filtering explicit toxicity, particularly when used in conjunction with more advanced annotation techniques.
3.2 Pre-processing